Many radioisotopes have important industrial, medical, and research applications. One of these is 60Co, which has a half-life of 5.20 years and decays by the emission of a beta particle (energy 0.310 MeV) and two gamma photons (energies 1.17 MeV and 1.33 MeV). A scientist wishes to prepare a 60Co sealed source that will have an activity of at least 10.8 Ci after 42.0 months of use.

a) If the activity is 10.8 Ci, how many 60Co atoms are in the source? Tries 0/10

b) What is the minimum number of nuclei in the source at the time of creation? Tries 0/10

c) What is the minimum initial mass of 60Co required?
Answer #1

Answer-

Given,

half life = 5.20 years or 1.64 * 108 sec   [1 year = 3.154 * 107 seconds]

Activity = 10.9 Ci or 4.033 * 1011 Bq [1 Ci = 3.7 * 1010 Bq]

Molar mass of 60Co = 59.933816 g/mol

time = 42 months or 1.104 * 108 sec [1 Month = 2.63 * 106 seconds]

a)

Number of atoms in 60 Co source = ?

We know that,

A = λ * N --------------------------3

where, A = activity in Bq

λ = decay constant

N = Number of Undecayed atoms in sample

Also,

half life = 0.693 / λ --------------4

From 3 and 4,

N = A / (0.693 /half life )

where, A = activity in Bq

N = Number of Undecayed atoms in sample

Put the values,

N = A / (0.693 /half life )

N = 4.033 * 1011 Bq / (0.693 /1.64 * 108 sec )

N = 9.54 * 1019 atoms

So, the number of atoms that will have activity of 10.8 Ci or 4.033 * 1011 Bq is, 9.54 * 1019 atoms. these are the atoms left after 42 months i.e. 1.104 * 108 sec [1 Month = 2.63 * 106 seconds]

Also,

Mass = Number of atoms/avogadro's number) * molar mass

mass in 9.54 * 1019 atoms = (9.54 * 1019 atoms / 6.022 * 1023) * 59.933816 g/mol

mass in 9.54 * 1019 atoms = 94.89 * 10-4 g

Also,

= 0.693 / half life

where, = decay rate

decay rate of 60 Co = 0.693/1.64 * 108

So, decay rate of 60 Co = 4.23 * 10-9 Bq

Also,

Nt = Ne-t

where, Nt = the amount of radioactive particles at time (t)

N0 = the amount of radioactive particles at time = 0

λ = rate of decay constant

t = time

Put the values,

Nt = Ne-t

94.89 * 10-4 g = (N) * e-0.00000000423 * 110400000

94.89 * 10-4 g = (N) * 0.627 [e-0.467 = 0.627]

N = 151.34 * 10-4 g

Number of atoms in 151.34 * 10-4 g of 60 Co is,

Number of atoms = (Mass/molar mass) * avogadro's number

Number of atoms = (151.34 * 10-4 g / 59.933816 g/mol) * 6.022 * 1023

Number of atoms = 1.521 * 1020 atoms

So, there are 1.521 * 1020 atoms present in the source due to which the activity of remaining atoms is 10.8 Ci after 42 months.

b)

Minimum Number of nuclie in the source = ?

The atoms needed to have Activity of 10.8 Ci after 42 months of use is 1.521 * 1020 atoms. So, the minimum number of atom or nuclei in source at the time of creation is same 1.521 * 1020 atoms.

c)

Number of atoms in the souce = 1.521 * 1020 atoms

Mass of atoms in the source = ?

We know that,

Mass = Number of atoms/avogadro's number) * molar mass

Put the values,

Mass = (1.521 * 1020 atoms / 6.022 * 1023 ) * 59.933816 g/mol

Mass = 1.514 * 10-2 g

So, the minimum mass in the source is 1.514 * 10-2 g